# Near-Optimal Very Low Rank Approximation at Sublinear Cost

### v16

@@ Appendix D Generation of abridged SRHT matrices
In this section we specify the family of abridged SRHT matrices, used in our tests for Algs. 2.1 and 4.1. They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [17, Sec. 3.1]). The $n\times n$ matrices $H_{n}$ are obtained in $t=\log_{2}(n)$ recursive steps, but we only perform $d\ll t$ steps, and the resulting abridged matrix $H_{d,d}$ can be multiplied by a vector by using $2dn$ additions and subtractions. SRHT matrices are obtained from the matrices $H_{n}$ by means of random sampling and scaling, which we also apply to the $d$-abridged Hadamard transform matrices $H_{d,d}$. They turn into $H_{n}$ for $d=t$ but are sparse fo $d\ll t$. Namely, we write $H_{d,0}:=I_{n/2^{d}}$ and then specify the following recursive process:
$H_{d,0}:=I_{n/2^{d}},~{}H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{}{\rm for}~{}i=0,1,\dots,d-1,$ (D.1)
@@
The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf. [11, Sec. 11]).

### v17

@@ Appendix C Generation of abridged SRHT matrices
In this section we specify the family of abridged SRHT matrices, used in our tests for Algs. 2.1 and 4.1. They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [14, Sec. 3.1]). The $n\times n$ matrices $H_{n}$ are obtained in $t=\log_{2}(n)$ recursive steps, but we only perform $d\ll t$ steps, and the resulting abridged matrix $H_{d,d}$ can be multiplied by a vector by using $2dn$ additions and subtractions. SRHT matrices are obtained from the matrices $H_{n}$ by means of random sampling and scaling, which we also apply to the $d$-abridged Hadamard transform matrices $H_{d,d}$. They turn into $H_{n}$ for $d=t$ but are sparse fo $d\ll t$. Namely, we write $H_{d,0}:=I_{n/2^{d}}$ and then specify the following recursive process:
$H_{d,0}:=I_{n/2^{d}},~{}H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{}{\rm for}~{}i=0,1,\dots,d-1,$ (C.1)
@@
The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf. [10, Sec. 11]).